UT Dec 01 2021 - Split time LUCI-AO Eng / AZ Science



Night Info

  • Sunset 17:19
  • Twilght End 18:40
  • Moonrise 03:49
  • Twilight Begins 05:37
  • Sunrise 06:59
  • Starting Config Bino LUCIs

Night Plan

  • ESM Testing 18:10-20:40hr
    • Run OSCO to verify stability of system in twilight
    • Prinary Goal is to verify operations of ESM Bin 4 (highest), Bin 1 and Bin 2 (secondary)
    • Observations will be run on a science field to verify the limits of Bin 4 fluxes and operation of the system
      • J2329-0522
    • Backup Bin 1 ESM testing
      • M52
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/M52/M52_R10pt7_ESM_SL.xml
  • Full AO Bin 2 testing
    • Use AO OT file: /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/AO/AO_Testing_V41.xml
      • RA 23hr-00hr
      • Star 5 (R=13.3)
      • Star 6 (R=13.7)
  • Return to AZ Science Queue
  • If conditions too poor for ESM - return to AZ queue - attempt ESM later in the night
    • Later target - M67
    • Bin 4 R=15.75
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/M67/Bin4/M67_Bin4_R15pt75_PA0_ESM_SL.xml
    • Bin 1 ESM tests for M67 if time
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/M67/Bin1
    • Mixed-mode scripts pre-generated

Summary

22:54 Juan Carlos is continuing tests for bin 4 for tonights operations. Bringing up the instruments.

modsDisp failed to open on ROBS so I have mods on obs2, LBC on robs and luci on obs5….

23:38 Field stop alignment checked, mods simSnap run. Running LBC Biases

0:20 David is opening

0:30 Yang came online and robs fixed its issue with the modsDisp all by itself. MODS and LBC are both on robs now.

0:35 David is executing pointing and collimation

0:48 Collimatin on a 10.3 magnitude star

0:50 OSCO target HD216399. Running in intervention mode. Optimized estimated magnitude is 9.6 magnitude.

Image is massively out of focus! Checking whether this is an AO focus or LUCI focus. GCS is in focus so this is a LUCI focus issue,

6:10. We took images witht he N30 & N375. We tried adjusted and stepping the detector step focus a few hundred steps but we still have a defocused image

1:12 stopAo and GCS guided image, taking image with N375. We get 2" seeing and out of focus images without AO

1.27 Cube was left in the beam, doug took it out. LUCI is in focus (dot circle on N30 in N375 in focus), so any out of focus was external to luci.

1:34 Preset to OSCO target, ACE active preset. 8.8 estimated magnitude on optimized OCAM image. There is a bite occuring around 10 O'clock. GOpt failed to stabilize within the first iteration, but successful the second time. Taking only one set with NCPA on and moving on.

luci1.202111201.0064-65 Skipping a whole bunch of frames, so reoptimizing.

1:44 Doug is executing an adsec stop start. Things are too funcky. The adsec is being a little pain in the butt and not powering up nicely.

1:52 Third time's the charm. Resending the OSCO preset. Seeing is 0.8" on the guider. Still have that bite/curled edge at 10 O'clock. NCPA on. Looks like we may have some overdriving
tt1.73, 0.30, 0.17 with the optical gain and NCPA. Looks nice, with pretty airy rings. Good range on the camera lens

luci1.202111201.0066-85
TN#015831
TN#020333

2:06 Sending ESM bin4 preset to J2329-0522. This is a target for AZ-2021B-007 that is being integrated into the ENG because of its limited opportunity and high priority for the AZ Q.

2:09 OCAM not live. Doug is interrupting command, changing binning on ocam to reset the ocam

2:12 Recovered and resuming center pupils. Camera lense is well within range. Estimated OCAM magnitude after applied gains 17.1, gains of 1.5, 0.05, 0. 1.1"

2:24 Doug drive up the h01 to 0.2 and the images look elongated, and the GCS image looks like a peanut. luci1.202111201.0088-90 are highly aberrated. Giving up.

2:29 Moving on to bin2 tests. Star5 bin2 PA0 test. G 172-41 preset sent. Seeing on GCS seeing 0.72"
Estimated magnitude on OCAM is 13.7, Optical Gain on, NCPA on, Gains 1.34, 0.12, 0.12.

luci.20211201.0092-93

TN#023923, TN#024035 - Strehl:53%

Good range on the camera lens. Corrections

TN#024339 luci2.20211201.0094

luci1.202111201.0086

2:50 Large offset, out of both guide and AO patrol field. Success.

2:54 Noted in luci1.20211201.99 the ocam briely showed not live, but returned in ~2-3 seconds.

3:03 Stop AO. We are reconfiguring to LBC's for AZ science.

Some ADI obs notes:

5:41 Sending collimation script for IP Tau.

OCAM estimated magnitude after optimization is 13.2. Preset magnitude was 12.46 in R.

We are at pixel 1042, 1152 to start (so 1.56” above center to start, slight offset from) Subwin SW 1 914 0 256 2048.

5:59 Science started

luci1.20211201.0116-

6:13 Seeing 1” on the DIMM, clear conditions. Stable and good correction on the AO, good range on the camera lens to start which should not change since we are not changing rotator positions.

7:58 WFS says closed but adsec says paused. Did not resume between image 203 and image 204. Doug is grabbing a few tracking numbers. AOS says “pausing” so pause command is still in the works and until that completes it can not resume.

8:17 Looks like the WFS arb lost communication. Restarted. The adsec executed a stopAO command when it should have just paused

8:26 We are back on source. Missed 5.5 degrees of field rotation. Pauses are now consistent between the adsec and wfs. We had to reoptimize, reapply gopt, and NCPA.

AOlog:

https://wiki.lbto.org/AdaptiveOptics/20211130_SX_Night


-- BarryRothberg - 29 Nov 2021
Topic revision: r5 - 01 Dec 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback